peter pays $85 for a watch after a discount of 20%. find the orginal price of the watch


1
Expert's answer
2021-02-24T06:03:29-0500

Let $x be the original price of the watch. Then $(x-85) is the sum of a discount.

So we obtain the equation:

"x - 85 = 0.2 x; \newline 
 x - 0.2x = 85; \newline
0.8x = 85; \newline
x =\frac{85}{0.8}; \newline
x = 106"

Thus, the original price of the watch is $106.